# Large Language Models in Argument Mining: A Survey

### v1

@@ 3.4.1 Key Point Analysis
Initial approaches relied on clustering or semantic similarity. For example, Bar-Haim et al.(Bar-Haim et al., 2020; 2021) introduced the idea of extracting key points and matching arguments to them using supervised and unsupervised approaches, which inspired multiple shared tasks. Phan et al.(Phan et al., 2021) proposed a simple yet highly accurate model using bi-encoders for key point matching, while Reimer et al. (Reimer et al., 2021) explored pretrained encoders and alignment strategies.
@@
With the rise of LLMs, researchers explored their generative and representational capabilities. Tang et al.(Tang et al., 2024b; a) examined aspect-based and prompted KPA for review summarization, highlighting the flexibility of LLMs in domain adaptation. Li et al.(Li et al., 2023) proposed an iterative clustering and abstractive summarization framework to automatically induce key points. More recently, Fukuma et al. (Fukuma et al., 2024) demonstrated that zero-shot KPA is feasible through LLM-based question generation and network analysis.
@@ 4.1 Prompt-based Argument Mining
Prompting has also been used to assess the sufficiency or quality of arguments. Li et al. (2025) explore prompt-based counterfactual calibration strategies to mitigate stance-related biases in LLMs. Chen & Eger (2025) use prompts for manipulation checks in evaluating the effect of emotional content on argument convincingness.
@@ 4.2 Instruction Tuning and Multi-task Fine-tuning
In the legal domain, instruction-tuned models such as ArgLegalSumm (Elaraby & Litman, 2022) improve summarization by aligning model outputs with argumentative structures. Elaraby et al. (2023) further enhanced legal summarization using re-ranking mechanisms driven by LLM-computed argument salience scores.
@@ 4.4 Synthetic Data and LLM-generated Supervision
This paradigm is not only cost-efficient but also enables the control and modulation of specific attributes such as diversity (van der Meer et al., 2024), exhaustiveness (Khosravani et al., 2024), and coherence (Shan & Lu, 2025)—factors crucial to high-quality argument outputs.
@@ 5.1 Convincingness Evaluation
Convincingness refers to how persuasive or compelling an argument is. Early studies such as Chalaguine & Schulz (2017) showed that simple linguistic and structural features could suffice for assessing convincingness in online debates. Later, more advanced models were introduced. Gu et al. (2018) explored incorporating topic aspects in evaluating comment convincingness on social platforms, while Potash et al. (2019) proposed ranking mechanisms for argument passages.
The impact of emotions has also been investigated. Chen & Eger (2025) introduced an LLM-based manipulation check system to explore whether emotional language biases human and machine assessments of argument convincingness.
@@ 5.2 Sufficiency and Quality
Further formal analysis of sufficiency was provided by Borg & Bex (2024), who explored the concepts of minimality, necessity, and sufficiency in the context of explanation and argumentation. Quality assessment tasks often intersect with sufficiency. For instance, Favreau et al. (2022) applied learning-to-rank BERT models to evaluate argument quality in educational settings.
@@ 5.3 Automatic Evaluation and End-to-End Frameworks
A novel coverage metric for summarization tasks was proposed by Khosravani et al. (2024), emphasizing exhaustiveness and accuracy in key point extraction, which reflects evaluation criteria in real-world persuasive writing. From a logic-oriented angle, Burrieza & Yuste-Ginel (2021) examined evaluation through justification logic in multi-agent environments. In education, Winkle et al. (2020) analyzed how teachers apply critical questions to evaluate student arguments, while Rach et al. (2020) benchmarked argument search and evaluation systems in dialogue settings.
@@ 5.4 Theoretical and Domain-Specific Perspectives
Biases in argument evaluation have also been studied. Diana et al. (2020) investigated adaptive instruction methods to mitigate value-based biases. Baccini & Hartmann (2022) formalized the well-known myside bias using Bayesian models.
@@ 6.1 Model Interpretability and Controllability
While LLMs show strong performance across tasks such as stance detection, argument generation, and summarization, their outputs often lack transparency and controllability. In high-stakes domains such as law (Elaraby & Litman, 2022; Xu & Ashley, 2023), healthcare, or political discourse, black-box behavior can hinder trust and adoption.
@@ 6.2 Data Quality, Bias, and Hallucination Risks
Several studies, such as Chen & Eger (2025) and Li et al. (2025), have raised concerns regarding the influence of emotion, ideology, or prompt phrasing on argument quality and stance. LLMs trained on massive web corpora may inherit or amplify biases, leading to unbalanced or misleading arguments.
@@ 6.5 Integration with Argumentation Theory
Bridging this gap may involve using theory-guided prompts, symbolic constraints, or structure-aware generation and evaluation systems (Borg & Bex, 2024; Palmieri, 2024). Such alignment could improve model reasoning, transparency, and downstream decision support.
@@ 6.6 Human-Centered and Interactive Argument Mining
Finally, the shift toward interactive systems—e.g., debate coaching bots, civic deliberation platforms, or educational feedback tools—demands models that support real-time, multi-turn, and user-aligned argument mining. This includes tracking dialog acts, modeling myside bias (Baccini & Hartmann, 2022), and adapting feedback to user values or learning goals (Diana et al., 2020).

### v2

@@ 3.4.1 Key Point Analysis
Initial approaches relied on clustering or semantic similarity. For example, Bar-Haim et al.(Bar-Haim et al., 2020, 2021) introduced the idea of extracting key points and matching arguments to them using supervised and unsupervised approaches, which inspired multiple shared tasks. Phan et al.(Phan et al., 2021) proposed a simple yet highly accurate model using bi-encoders for key point matching, while Reimer et al. (Reimer et al., 2021) explored pretrained encoders and alignment strategies.
@@
With the rise of LLMs, researchers explored their generative and representational capabilities. Tang et al.(Tang et al., 2024b, a) examined aspect-based and prompted KPA for review summarization, highlighting the flexibility of LLMs in domain adaptation. Li et al.(Li et al., 2023) proposed an iterative clustering and abstractive summarization framework to automatically induce key points. More recently, Fukuma et al. (Fukuma et al., 2024) demonstrated that zero-shot KPA is feasible through LLM-based question generation and network analysis.
@@ 4.1 Prompt-based Argument Mining
Prompting has also been used to assess the sufficiency or quality of arguments. Li et al. (2025) explore prompt-based counterfactual calibration strategies to mitigate stance-related biases in LLMs. Chen and Eger (2025) use prompts for manipulation checks in evaluating the effect of emotional content on argument convincingness.
@@ 4.2 Instruction Tuning and Multi-task Fine-tuning
In the legal domain, instruction-tuned models such as ArgLegalSumm (Elaraby and Litman, 2022) improve summarization by aligning model outputs with argumentative structures. Elaraby et al. (2023) further enhanced legal summarization using re-ranking mechanisms driven by LLM-computed argument salience scores.
@@ 4.4 Synthetic Data and LLM-generated Supervision
This paradigm is not only cost-efficient but also enables the control and modulation of specific attributes such as diversity (van der Meer et al., 2024), exhaustiveness (Khosravani et al., 2024), and coherence (Shan and Lu, 2025)—factors crucial to high-quality argument outputs.
@@ 5.1 Convincingness Evaluation
Convincingness refers to how persuasive or compelling an argument is. Early studies such as Chalaguine and Schulz (2017) showed that simple linguistic and structural features could suffice for assessing convincingness in online debates. Later, more advanced models were introduced. Gu et al. (2018) explored incorporating topic aspects in evaluating comment convincingness on social platforms, while Potash et al. (2019) proposed ranking mechanisms for argument passages.
The impact of emotions has also been investigated. Chen and Eger (2025) introduced an LLM-based manipulation check system to explore whether emotional language biases human and machine assessments of argument convincingness.
@@ 5.2 Sufficiency and Quality
Further formal analysis of sufficiency was provided by Borg and Bex (2024), who explored the concepts of minimality, necessity, and sufficiency in the context of explanation and argumentation. Quality assessment tasks often intersect with sufficiency. For instance, Favreau et al. (2022) applied learning-to-rank BERT models to evaluate argument quality in educational settings.
@@ 5.3 Automatic Evaluation and End-to-End Frameworks
A novel coverage metric for summarization tasks was proposed by Khosravani et al. (2024), emphasizing exhaustiveness and accuracy in key point extraction, which reflects evaluation criteria in real-world persuasive writing. From a logic-oriented angle, Burrieza and Yuste-Ginel (2021) examined evaluation through justification logic in multi-agent environments. In education, Winkle et al. (2020) analyzed how teachers apply critical questions to evaluate student arguments, while Rach et al. (2020) benchmarked argument search and evaluation systems in dialogue settings.
@@ 5.4 Theoretical and Domain-Specific Perspectives
Biases in argument evaluation have also been studied. Diana et al. (2020) investigated adaptive instruction methods to mitigate value-based biases. Baccini and Hartmann (2022) formalized the well-known myside bias using Bayesian models.
@@ 6.1 Model Interpretability and Controllability
While LLMs show strong performance across tasks such as stance detection, argument generation, and summarization, their outputs often lack transparency and controllability. In high-stakes domains such as law (Elaraby and Litman, 2022; Xu and Ashley, 2023), healthcare, or political discourse, black-box behavior can hinder trust and adoption.
@@ 6.2 Data Quality, Bias, and Hallucination Risks
Several studies, such as Chen and Eger (2025) and Li et al. (2025), have raised concerns regarding the influence of emotion, ideology, or prompt phrasing on argument quality and stance. LLMs trained on massive web corpora may inherit or amplify biases, leading to unbalanced or misleading arguments.
@@ 6.5 Integration with Argumentation Theory
Bridging this gap may involve using theory-guided prompts, symbolic constraints, or structure-aware generation and evaluation systems (Borg and Bex, 2024; Palmieri, 2024). Such alignment could improve model reasoning, transparency, and downstream decision support.
@@ 6.6 Human-Centered and Interactive Argument Mining
Finally, the shift toward interactive systems—e.g., debate coaching bots, civic deliberation platforms, or educational feedback tools—demands models that support real-time, multi-turn, and user-aligned argument mining. This includes tracking dialog acts, modeling myside bias (Baccini and Hartmann, 2022), and adapting feedback to user values or learning goals (Diana et al., 2020).